# Global survey of star clusters in the Milky Way VIII. Cluster formation and evolution

### v1

@@ Abstract
Methods. The model for cluster evolution is comprised of a two-section cluster initial mass function, constant cluster formation rate, and a mass loss function. This mass loss function represents a supervirial phase after sudden expulsion of the remaining gas, cluster mass loss due to stellar evolution and gradual cluster dissolution driven by internal dynamics and the Galactic tidal field. We construct different estimators of model fitness based on $\chi^{2}$-statistics, the Kullback-Leibler divergence (KLD) and a maximum-likelihood approach, taking into account the uncertainty of our observed cluster parameters. Using these estimators and Markov Chain Monte Carlo (MCMC) sampling, we obtain best-fit values and posterior distributions for a selection of model parameters.
@@ 1 Introduction
Recently, Just et al. (2023) derived tidal masses for open clusters of the MWSC catalogue (see Kharchenko et al. 2012, 2013; Schmeja et al. 2014; Scholz et al. 2015) and constructed CMFs for different cluster age ranges for a completeness-corrected sample of Galactic open clusters. They described the CMFs in terms of a low-mass and a high-mass power-law slope, and found that while the high-mass slope remains close to one for all ages, the low-mass slope is initially shallow for clusters younger than 20 Myr. For older clusters, the low-mass slope becomes approximately $-0.7$ and changes little after the first 20 Myr. As the cluster ages increase, the breakpoint of this low-mass slope is shifted to higher masses. They presented a simple analytical model for cluster formation and evolution able to reproduce the general shape and several features of these CMFs, in particular the high-mass and low-mass slopes for ages above 20 Myr, as well as the shift of the CMFs towards higher masses with increasing age.
It is our aim to investigate this model now in the context of the full 2D CAMF, which describes the Milky Way cluster surface density as a function in the cluster age-mass plane. We will construct the empirical CAMF and provide a quantitative analysis of the model’s correspondence to observations. Further, we will perform a study of the model parameter space in order to find better estimates of the model parameters and related quantities of interest in order to improve the quality of the model as a description of the observed cluster population.
@@ 2 Data
In contrast to cluster ages, data on cluster masses for Milky Way clusters is not easy to compute and therefore not very abundant. One can mention the earlier data of Lada & Lada (2003) based on a collection of data of 76 embedded clusters; of Lamers et al. (2005) and of Piskunov et al. (2007) for 520 clusters from early release of the Catalog of Open Cluster Data (COCD, Kharchenko et al. 2005), and of Piskunov et al. (2008) for 650 clusters from the final release of COCD. Later Kharchenko et al. (2013) based on MWSC data determined tidal radii and masses (Just et al. 2023) for 3061 open clusters. The recent releases of the Gaia survey (Gaia Collaboration et al. 2016b) allowed for new determinations of cluster masses based on the DR2 and (E)DR3 releases: Meingast et al. (2021) for ten nearby clusters, Cordoni et al. (2023) for 78 clusters, Almeida et al. (2023) for 773 objects, and by Hunt & Reffert (2024) for 6956 clusters previously identified by them.
@@
The measurement of tidal radii and computation of the corresponding bound masses using a model of the Galactic gravitational potential (Piskunov et al. 2008; Kharchenko et al. 2013) is an alternative which is free from the above bias, and is probably the only direct measure of cluster mass available in practice. This has only become possible with the publication of all-sky surveys such as Hipparcos/Tycho (ESA 1997), the Two Micron All Sky Survey (2MASS, Skrutskie et al. 2006), the PPMXL Catalog of Positions and Proper Motions on the ICRS (PPMXL, Röser et al. 2010), and Gaia (e.g. Gaia Collaboration et al. 2018, 2021, 2023). Their use has opened the possibility to determine extensive lists of cluster tidal radii and masses, which can be used as an independent source of data for the CMF. It should be noted, however, that the determination of tidal masses is not without its own difficulties; such as difficulties with fixing the cluster outer boundary in dense or variable extinction environments. However, the independence from the age determination and weaker assumptions about the behaviour of the unseen fraction of the cluster population compensate for this drawback.
Once the cluster masses are fixed, one can construct the CMF, or its 2D analog, the age-mass distribution. The main requirement is full statistical representativeness of the cluster sample over the entire mass range. This means that the sample should provide a sufficient number of objects even at the extremes of the mass/age scales. In practice, a sample of several hundred clusters at the least is needed to resolve interesting features in the age-mass plane. Further, the application of a proper data completeness model to the raw distributions is necessary to account for varying completeness at different masses/ages. These conditions cut off all the poorly populated samples mentioned above. The Hunt & Reffert (2024) sample represents the most populated list of cluster masses and ages ever published. However, the cluster age and mass distributions derived from these data are based only on a provisional, mass-limited completeness model. The magnitude-limited nature of the Gaia survey suggests a completeness model based on cluster luminosities, for which a completeness distance determined solely by cluster masses is only a very limited substitute. However, cluster luminosities are neither constructed nor discussed by Hunt & Reffert (2024), and no follow-up studies of the catalogue’s completeness have been published at the time of writing.
@@
The MWSC catalogue (Kharchenko et al. 2012, 2013), represents a heliocentric sample of Galactic star clusters built using astrometry of PPMXL (Röser et al. 2010) in combination with near-infrared photometry from 2MASS (Skrutskie et al. 2006). It contains basic parameters for 3061 open clusters (Kharchenko et al. 2013), 202 of which were newly discovered as part of the MWSC survey (Schmeja et al. 2014; Scholz et al. 2015). Investigation by Kharchenko et al. (2016) determined that the MWSC sample is magnitude-limited. Just et al. (2023) extended the list of basic cluster parameters by the cluster tidal masses.
As the MWSC sample of open clusters can be characterised as magnitude-limited, it is necessary to address the issue of in-sample bias due to variable completeness. For this purpose, we follow here the approach of ‘magnitude-dependent completeness limits’ used by Piskunov et al. (2018); Just et al. (2023), which requires the assumption of a uniform distribution of clusters in the Galactic plane in the solar neighbourhood. Each cluster is assigned an individual completeness distance
@@
depending on its integrated magnitude $I(M_{K_{\mathrm{S}}})$ in the $K_{\mathrm{S}}$ passband. Here we adopt the values $p=0.36\,\mathrm{kpc}$ and $q=0.54\,\mathrm{kpc\,mag^{-1}}$ from Just et al. (2023). Only clusters which have a heliocentric distance projected to the Galactic plane $d_{xy}$ smaller than their individual magnitude-dependent completeness limit, that is, which fall inside their completeness area
@@
We consider now the distribution of clusters in the plane spanned by cluster ages and masses.
For ease and brevity of notation, we denote in the following
@@
The $P_{i,k,l}$ are computed by assuming a Gaussian error in logarithmic age-mass space. We define the Gaussian function
@@
Uncertainties in the tidal masses are given in the MWSC catalogue for all clusters for which tidal masses were determined. However, only a small subset of clusters has explicitly given age uncertainties. We refer to Kharchenko et al. (2013, Table 6) for upper and lower bounds on the uncertainty of cluster age determination and use estimated uncertainties of $\delta_{\tau}=0.1$ for clusters with age $\tau>8.2$ and $\delta_{\tau}=0.14$ for clusters with age $\tau\leq 8.2$.
We note that this approach is conceptually very similar to that of kernel density estimation, where some kernel function is used to estimate an underlying smooth density function from a set of discrete data points. Here, however, kernels of different size are used for the different clusters, and directly represent the uncertainties of cluster age-mass determination.
@@
Our completeness-corrected sample of MWSC clusters spans a range of tidal masses $\mu=0$–4.4 and ages $\tau=6$–9.8. For the sake of our analysis, we exclude 57 young clusters with $\tau<6.5$. The reasoning for this is twofold: Firstly, the age determination was done using Padova isochrones, for which the lower age limit is $\tau=6.6$, and secondly clusters younger than 3 Myr may still be in the gas-embedded, star-forming phase, which we do not consider here. This leaves us with a sample of 2170 clusters.
@@
In Fig. 1 the CAMFs constructed according to Eqs. (6) and (7) are plotted for logarithmic step sizes $\Delta\tau=0.1$ and $\Delta\mu=0.1$.
We can see already some features of the CAMF that bear relevance for modelling.
@@ 3 Cluster formation and evolution model
We use a model for the CAMF that is based directly on the model presented in Just et al. (2023), which itself is an extension of an earlier model used in Piskunov et al. (2018) to model the cluster age distribution.
@@
via the product of the cluster formation rate (CFR) $\Psi(t)$ and the cluster initial mass function (CIMF) $f(M)$, where $M$ is the initial cluster mass directly after formation.
The CFR gives the rate at which clusters form as a function of time in terms of cluster surface density per unit time and the CIMF gives the initial mass distribution of newly formed clusters, here normalised to unity.
The initial mass $M$ of a cluster and its current mass $m$ at time $t$ are related by a bound-mass function $m(M,t)$ which models cluster mass evolution and is implicitly assumed to be invertible, such that at a given age, exactly one initial mass corresponds to any given $m$.
As the mass evolution of real clusters is a deeply stochastic process (driven for instance by IMF sampling, stellar evolution and complex gravitational dynamics; cf. e.g. Wang & Jerabkova 2021; Ernst et al. 2008) which depends not just on the initial mass but on the specific dynamical state, this model of mass evolution must be understood on the population level as the average mass evolution of an ensemble of individual clusters with fixed initial mass $M$.
This mass evolution model also accounts for the destruction of clusters, as clusters with an evolved bound mass of 0 have become fully unbound and dissolved, and no longer contribute to the CAMF.
@@ 3.1 Cluster formation rate
The CFR $\Psi(t)$ gives the number of newly formed clusters per squared kpc of Galactic disc surface and per Myr.
We use a constant CFR
@@
This parameter $\beta$ was chosen such that the modelled total cluster surface density $\Sigma$ was equal to the observed total surface density $\sum\limits_{i=1}^{N_{\mathrm{Cl}}}S_{i}^{-1}$. The physical intuition behind this is that the total cluster content corresponds to the CFR by integration over time, taking the dissolution of clusters into account via their mass evolution.
The total cluster surface density
@@
is computed by integrating the cluster age-mass distribution for integration bounds $\log t/\mathrm{yr}=6.5$–10 and $\log m/\mathrm{M}_{\odot}=0$–4.5 corresponding to the observed age-mass range of open star clusters in our sample.
@@ 3.2 Cluster initial mass function
The CIMF gives the mass distribution of clusters at initial age $t=0$.
We use a CIMF that is informed by the shape of the CMF for young clusters with ages up to 20 Myr (as in Just et al. 2023, Fig. 9). It takes the shape of a two-slope broken power law with a smooth transition region and allows for an exponential, Schechter-type cut-off. Our CIMF, defined as the fraction of clusters $\mathrm{d}N$ in initial mass interval $\mathrm{d}M$, is given by
@@
with low-mass power-law index $x_{1}$ for $M<M_{\ast}$ and high-mass power-law index $x_{2}$ for $M>M_{\ast}$. The exponent $s$ controls the ‘sharpness’ of the transition between the power laws. A larger value of $s$ results in a less extended transition region around $M_{\ast}$ and vice versa for lower values of $s$. Finally, $m_{\mathrm{S}}$ is the Schechter cut-off mass above which the CIMF drops exponentially.
The CIMF is normalised to unity by the normalisation constant $k$ on the initial mass interval with lower limit $m_{\mathrm{lower}}=2\,\mathrm{M}_{\odot}$ and no upper limit.
@@ 3.3 Cluster bound-mass function
where $n_{\mathrm{L}}(t)$ is the mass loss factor from stellar evolution, $n_{\mathrm{V}}(t)$ gives the bound mass fraction from violent relaxation, and $n_{\mathrm{sec}}(M,t)$ is the bound mass fraction from secular dynamical evolution. The mass loss factor from stellar evolution depends on the details of stellar evolution as well as the stellar initial mass function. Following Lamers et al. (2010), we approximate the remaining stellar mass fraction $n_{\mathrm{L}}(t)$ by a third order polynomial in logarithmic time
$n_{\mathrm{L}}(t)=\sum\limits_{i=0}^{3}p_{i}\left[\log(t/\mathrm{Myr})\right]^{i}.$ (21)
@@
The violent relaxation phase that occurs in infant clusters after gas expulsion has been investigated in detail by Shukirgaliyev et al. (2017, 2018, 2019, 2021) using N-body models. They found that the fraction of stars $n_{\mathrm{b}}$ remaining bound to a cluster after virialisation depends strongly on the global star formation efficiency, but not on the initial mass and weakly on the initial size of the cluster. This motivates us to formulate the bound mass fraction from violent relaxation $n_{\mathrm{V}}(t)$ as a function of time only, without dependence on the initial cluster mass $M$. We define
@@
We note that some star clusters do not survive gas expulsion and rapidly dissolve in this supervirial phase. Cluster survivability is linked to star forming efficiency (Shukirgaliyev et al. 2017, 2018), density profile (Shukirgaliyev et al. 2021), primordial mass segregation (Vesperini et al. 2009; Haghi et al. 2014; Brinkmann et al. 2017), initial stellar mass function (Haghi et al. 2020), and crucially the dynamics of gas expulsion itself (Smith et al. 2013; Zamora-Avilés et al. 2019).
We do not model this early rapid disruption of clusters due to the low number of clusters observed at young ages, which makes estimating the fraction of clusters affected difficult and limits the impact on the CAMF.
The final factor $n_{\mathrm{sec}}(M,t)$ gives the bound mass fraction from secular evolution of a star cluster in the Galactic tidal field. This mass loss characterised by slow evaporation and eventual dissolution must depend on the initial cluster mass in order to recover a dependence of the cluster lifetime on the initial cluster mass.
Our approach here mirrors that of Lamers & Gieles (2006); Lamers et al. (2005), with both the mass loss rate and the cluster lifetime being characterised by power laws in initial mass. However, we do not consider different effects separately and thus do not fix parameters a priori.
We define the secular bound fraction as
@@
We use the lifetime-mass relation
@@
This lifetime follows a broken two-slope power-law with index $1-a_{1}$ for low masses $M\ll M_{\mathrm{br}}$ and index $1-a_{1}-a_{2}$ for high masses $M\gg M_{\mathrm{br}}$. The constant $c$ governs the overall rate of secular mass loss. We note that this lifetime is parametrised slightly differently than in Just et al. (2023). There, $c$ was scaled to the lifetime at $M_{\mathrm{br}}$, while $c$ is here scaled to the lifetimes of low-mass clusters $M\ll M_{\mathrm{br}}$. As such, the value of $c$ given in Table 1 differs from the one of Just et al. (2023) by a factor of $\left(M_{\mathrm{br}}/\mathrm{M}_{\odot}\right)^{1-a_{1}}$.

### v2

@@ Abstract
Methods. The model for cluster evolution is comprised of a two-section cluster initial mass function, constant cluster formation rate, and a mass loss function. This mass loss function represents a supervirial phase after sudden expulsion of the remaining gas, cluster mass loss due to stellar evolution and gradual cluster dissolution driven by internal dynamics and the Galactic tidal field. We constructed different estimators of model fitness based on $\chi^{2}$-statistics, the Kullback-Leibler divergence (KLD) and a maximum-likelihood approach, taking into account the uncertainty of our observed cluster parameters. Using these estimators and Markov chain Monte Carlo (MCMC) sampling, we obtained best-fit values and posterior distributions for a selection of model parameters.
@@ 1 Introduction
Recently, Just et al. (2023) derived tidal masses for open clusters of the MWSC catalogue (see Kharchenko et al. 2012, 2013; Schmeja et al. 2014; Scholz et al. 2015) and constructed CMFs for different cluster age ranges for a completeness-corrected sample of Galactic open clusters. They described the CMFs in terms of a low-mass and a high-mass power-law slope, and find that while the high-mass slope remains close to one for all ages, the low-mass slope is initially shallow for clusters younger than 20 Myr. For older clusters, the low-mass slope becomes approximately $-0.7$ and changes little after the first 20 Myr. As the cluster ages increase, the breakpoint of this low-mass slope is shifted to higher masses. They present a simple analytical model for cluster formation and evolution able to reproduce the general shape and several features of these CMFs, in particular the high-mass and low-mass slopes for ages above 20 Myr, as well as the shift of the CMFs towards higher masses with increasing age.
It is our aim to investigate this model now in the context of the full 2D CAMF, which describes the Milky Way cluster surface density as a function in the cluster age-mass plane. We construct the empirical CAMF and provide a quantitative analysis of the model’s correspondence to observations. Further, we perform a study of the model parameter space in order to find better estimates of the model parameters and related quantities of interest in order to improve the quality of the model as a description of the observed cluster population.
@@ 2 Data
In contrast to cluster ages, data on cluster masses for Milky Way clusters is not easy to compute and therefore not very abundant. One can mention the earlier data of Lada & Lada (2003) based on a collection of data of 76 embedded clusters; of Lamers et al. (2005) and of Piskunov et al. (2007) for 520 clusters from early release of the Catalog of Open Cluster Data (COCD, Kharchenko et al. 2005), and of Piskunov et al. (2008) for 650 clusters from the final release of COCD. Later Kharchenko et al. (2013) based on MWSC data determined tidal radii and masses (Just et al. 2023) for 3061 open clusters. The recent releases of the Gaia survey (Gaia Collaboration et al. 2016b) allow for new determinations of cluster masses based on the DR2 and (E)DR3 releases: Meingast et al. (2021) for ten nearby clusters, Cordoni et al. (2023) for 78 clusters, Almeida et al. (2023) for 773 objects, and by Hunt & Reffert (2024) for 6956 clusters previously identified by them.
@@
The measurement of tidal radii and computation of the corresponding bound masses using a model of the Galactic gravitational potential (Piskunov et al. 2008; Kharchenko et al. 2013) is an alternative which is free from the above bias, and is probably the only direct measure of cluster mass available in practice. This has only become possible with the publication of all-sky surveys such as Hipparcos/Tycho (ESA 1997), the Two Micron All Sky Survey (2MASS, Skrutskie et al. 2006), the PPMXL Catalog of Positions and Proper Motions on the ICRS (PPMXL, Röser et al. 2010), and Gaia (e.g. Gaia Collaboration et al. 2018, 2021, 2023). Their use opens the possibility to determine extensive lists of cluster tidal radii and masses, which can be used as an independent source of data for the CMF. It should be noted, however, that the determination of tidal masses is not without its own difficulties; such as difficulties with fixing the cluster outer boundary in dense or variable extinction environments. However, the independence from the age determination and weaker assumptions about the behaviour of the unseen fraction of the cluster population compensate for this drawback.
Once the cluster masses are fixed, one can construct the CMF, or its 2D analogue, the age-mass distribution. The main requirement is full statistical representativeness of the cluster sample over the entire mass range. This means that the sample should provide a sufficient number of objects even at the extremes of the mass and age scales. In practice, a sample of several hundred clusters at the least is needed to resolve interesting features in the age-mass plane. Further, the application of a proper data completeness model to the raw distributions is necessary to account for varying completeness at different masses and ages. These conditions cut off all the poorly populated samples mentioned above. The Hunt & Reffert (2024) sample represents the most populated list of cluster masses and ages ever published. However, the cluster age and mass distributions derived from these data are based only on a provisional, mass-limited completeness model. The magnitude-limited nature of the Gaia survey suggests a completeness model based on cluster luminosities, for which a completeness distance determined solely by cluster masses is only a very limited substitute. However, cluster luminosities are neither constructed nor discussed by Hunt & Reffert (2024), and no follow-up studies of the catalogue’s completeness have been published at the time of writing.
@@
The MWSC catalogue (Kharchenko et al. 2012, 2013), represents a heliocentric sample of Galactic star clusters built using astrometry of PPMXL (Röser et al. 2010) in combination with near-infrared photometry from 2MASS (Skrutskie et al. 2006). It contains basic parameters for 3061 open clusters (Kharchenko et al. 2013), 202 of which are newly discovered as part of the MWSC survey (Schmeja et al. 2014; Scholz et al. 2015). Investigation by Kharchenko et al. (2016) determines that the MWSC sample is magnitude-limited. Just et al. (2023) extend the list of basic cluster parameters by the cluster tidal masses.
As the MWSC sample of open clusters can be characterised as magnitude-limited, it is necessary to address the issue of in-sample bias due to variable completeness. For this purpose, we followed here the approach of ‘magnitude-dependent completeness limits’ used by Piskunov et al. (2018); Just et al. (2023), which requires the assumption of a uniform distribution of clusters in the Galactic plane in the solar neighbourhood. Each cluster was assigned an individual completeness distance
@@
depending on its integrated magnitude $I(M_{K_{\mathrm{S}}})$ in the $K_{\mathrm{S}}$ passband. Here we adopted the values $p=0.36\,\mathrm{kpc}$ and $q=0.54\,\mathrm{kpc\,mag^{-1}}$ from Just et al. (2023). Only clusters which have a heliocentric distance projected to the Galactic plane $d_{xy}$ smaller than their individual magnitude-dependent completeness limit, that is, which fall inside their completeness area
@@
We consider now the distribution of clusters in the plane spanned by cluster ages and masses. For ease and brevity of notation, we denote in the following
@@
The $P_{i,k,l}$ were computed by assuming a Gaussian error in logarithmic age-mass space. We define the Gaussian function
@@
Uncertainties in the tidal masses are given in the MWSC catalogue for all clusters for which tidal masses were determined. However, only a small subset of clusters has explicitly given age uncertainties. We referred to Kharchenko et al. (2013, Table 6) for upper and lower bounds on the uncertainty of cluster age determination and used estimated uncertainties of $\delta_{\tau}=0.1$ for clusters with age $\tau>8.2$ and $\delta_{\tau}=0.14$ for clusters with age $\tau\leq 8.2$.
We note that this approach is conceptually very similar to that of kernel density estimation, where some kernel function is used to estimate an underlying smooth density function from a set of discrete data points. Here, however, kernels of different size were used for the different clusters, and directly represented the uncertainties of cluster age-mass determination.
@@
Our completeness-corrected sample of MWSC clusters spans a range of tidal masses $\mu=0$–4.4 and ages $\tau=6$–9.8. For the sake of our analysis, we excluded 57 young clusters with $\tau<6.5$. The reasoning for this is twofold: Firstly, the age determination was done using Padova isochrones, for which the lower age limit is $\tau=6.6$, and secondly clusters younger than 3 Myr may still be in the gas-embedded, star-forming phase, which we do not consider here. This leaves us with a sample of 2170 clusters.
@@
In Fig. 1 the CAMFs constructed according to Eqs. (6) and (7) are plotted for logarithmic step sizes $\Delta\tau=0.1$ and $\Delta\mu=0.1$. We can see already some features of the CAMF that bear relevance for modelling.
@@ 3 Cluster formation and evolution model
We used a model for the CAMF that is based directly on the model presented in Just et al. (2023), which itself is an extension of an earlier model used in Piskunov et al. (2018) to model the cluster age distribution.
@@
via the product of the cluster formation rate (CFR) $\Psi(t)$ and the cluster initial mass function (CIMF) $f(M)$, where $M$ is the initial cluster mass directly after formation. The CFR gives the rate at which clusters form as a function of time in terms of cluster surface density per unit time and the CIMF gives the initial mass distribution of newly formed clusters, here normalised to unity.
The initial mass $M$ of a cluster and its current mass $m$ at time $t$ are related by a bound-mass function $m(M,t)$ which models cluster mass evolution and is implicitly assumed to be invertible, such that at a given age, exactly one initial mass corresponds to any given $m$. The mass evolution of real clusters is a deeply stochastic process (driven for instance by IMF sampling, stellar evolution and complex gravitational dynamics; cf. e.g. Wang & Jerabkova 2021; Ernst et al. 2008) which depends not just on the initial mass but on the specific dynamical state. Therefore, this model of mass evolution must be understood on the population level as the average mass evolution of an ensemble of individual clusters with fixed initial mass $M$. This mass evolution model also accounts for the destruction of clusters, as clusters with an evolved bound mass of 0 have become fully unbound and dissolved, and no longer contribute to the CAMF.
@@ 3.1 Cluster formation rate
The CFR $\Psi(t)$ gives the number of newly formed clusters per squared kpc of Galactic disc surface and per Myr. We used a constant CFR
@@
This parameter $\beta$ was chosen such that the modelled total cluster surface density $\Sigma$ was equal to the observed total surface density $\sum\limits_{i=1}^{N_{\mathrm{Cl}}}S_{i}^{-1}$. The physical intuition behind this is that the total cluster content corresponds to the CFR by integration over time, taking the dissolution of clusters into account via their mass evolution. The total cluster surface density
@@
was computed by integrating the cluster age-mass distribution for integration bounds $\log t/\mathrm{yr}=6.5$–10 and $\log m/\mathrm{M}_{\odot}=0$–4.5 corresponding to the observed age-mass range of open star clusters in our sample.
@@ 3.2 Cluster initial mass function
The CIMF gives the mass distribution of clusters at initial age $t=0$. We used a CIMF that is informed by the shape of the CMF for young clusters with ages up to 20 Myr (as in Just et al. 2023, Fig. 9). It takes the shape of a two-slope broken power law with a smooth transition region and allows for an exponential, Schechter-type cut-off. Our CIMF, defined as the fraction of clusters $\mathrm{d}N$ in initial mass interval $\mathrm{d}M$, is given by
@@
with low-mass power-law index $x_{1}$ for $M<M_{\ast}$ and high-mass power-law index $x_{2}$ for $M>M_{\ast}$. The exponent $s$ controls the ‘sharpness’ of the transition between the power laws. A larger value of $s$ results in a less extended transition region around $M_{\ast}$ and vice versa for lower values of $s$. Finally, $m_{\mathrm{S}}$ is the Schechter cut-off mass above which the CIMF drops exponentially. The CIMF was normalised to unity by the normalisation constant $k$ on the initial mass interval with lower limit $m_{\mathrm{lower}}=2\,\mathrm{M}_{\odot}$ and no upper limit.
@@ 3.3 Cluster bound-mass function
where $n_{\mathrm{L}}(t)$ is the mass loss factor from stellar evolution, $n_{\mathrm{V}}(t)$ gives the bound mass fraction from violent relaxation, and $n_{\mathrm{sec}}(M,t)$ is the bound mass fraction from secular dynamical evolution.
The mass loss factor from stellar evolution depends on the details of stellar evolution as well as the stellar initial mass function. Following Lamers et al. (2010), we approximated the remaining stellar mass fraction $n_{\mathrm{L}}(t)$ by a third order polynomial in logarithmic time
$n_{\mathrm{L}}(t)=\sum\limits_{i=0}^{3}p_{i}\left[\log(t/\mathrm{Myr})\right]^{i},$ (21)
@@
The violent relaxation phase that occurs in infant clusters after gas expulsion has been investigated in detail by Shukirgaliyev et al. (2017, 2018, 2019, 2021) using N-body models. They find that the fraction of stars $n_{\mathrm{b}}$ remaining bound to a cluster after virialisation depends strongly on the global star formation efficiency, but not on the initial mass and weakly on the initial size of the cluster. This motivates us to formulate the bound mass fraction from violent relaxation $n_{\mathrm{V}}(t)$ as a function of time only, without dependence on the initial cluster mass $M$. We defined
@@
We note that some star clusters do not survive gas expulsion and rapidly dissolve in this supervirial phase. Cluster survivability is linked to star forming efficiency (Shukirgaliyev et al. 2017, 2018), density profile (Shukirgaliyev et al. 2021), primordial mass segregation (Vesperini et al. 2009; Haghi et al. 2014; Brinkmann et al. 2017), initial stellar mass function (Haghi et al. 2020), and crucially the dynamics of gas expulsion itself (Smith et al. 2013; Zamora-Avilés et al. 2019). We did not model this early rapid disruption of clusters due to the low number of clusters observed at young ages, which makes estimating the fraction of clusters affected difficult and limits the impact on the CAMF.
The final factor $n_{\mathrm{sec}}(M,t)$ gives the bound mass fraction from secular evolution of a star cluster in the Galactic tidal field. This mass loss characterised by slow evaporation and eventual dissolution must depend on the initial cluster mass in order to recover a dependence of the cluster lifetime on the initial cluster mass. Our approach here mirrors that of Lamers & Gieles (2006); Lamers et al. (2005), with both the mass loss rate and the cluster lifetime being characterised by power laws in initial mass. However, we did not consider different effects separately and thus did not fix parameters a priori.
We defined the secular bound fraction as
@@
We used the lifetime-mass relation
@@
This lifetime follows a broken two-slope power-law with index $1-a_{1}$ for low masses $M\ll M_{\mathrm{br}}$ and index $1-a_{1}-a_{2}$ for high masses $M\gg M_{\mathrm{br}}$. The constant $c$ governs the overall rate of secular mass loss. We note that this lifetime is parametrised slightly differently than in Just et al. (2023). There, $c$ was scaled to the lifetime at $M_{\mathrm{br}}$, while $c$ was here scaled to the lifetimes of low-mass clusters $M\ll M_{\mathrm{br}}$. As such, the value of $c$ given in Table 1 differs from the one of Just et al. (2023) by a factor of $\left(M_{\mathrm{br}}/\mathrm{M}_{\odot}\right)^{1-a_{1}}$.